From a 2,000 μg/mL BSA stock, create 1 mL of each of the following stock solutions in 0.02 M phosphate buffer using individual microcentrifuge tubes: 50 μg/mL, 250 μg/mL, 500 μg/mL, 1,000 μg/mL, 1,250 μg/mL, 1,500 μg/mL. Calculate the volume of BSA stock that will be required to make the standard solutions needed to create the BSA standard curve. Be sure to show your work and include the volume of 0.02 M phosphate buffer required to reach a final volume of 1 mL.
Answer #1

Okay before I start with the answer, let me explain something about the formula which is used to determine the amount of stock solution we need to make a standard solution.

The formula is M1 * V1 = M2 * V2

In this formula,

M1= The concentration (molarity) of the stock solution

V1= The amount (volume) which we have to take from the stock solution to make the standard solution.

M2= It is the concentration (molarity) of the standard solution which we have to make.

V2= It is the final volume of our standard solution which we are making.

So let's take one standard solution from your question.

You have to make 50 μg/ml of stock from the 2000 μg/ml. You've to make 1 ml of this 50 μg/ml solution.

So here:-

M1 is 2000  μg/ml

V1 is what we have to find.

M2 is 50  μg/ml

V2 is 1 ml ( we will convert it into μl by multiplying quantity in ml with 1000, thus 1 ml = 1000 μl)

So putting the value in the formula, we have

2000 * V1 = 50 * 1000

2000 * V1 = 50,000 (Multiplying 50 with 1000)

V1 = 50,000 / 2000 (Since we took over 2000 to other side of the equation, the signs have reversed)

V1 = 25 μl (since it's volume, the unit is litre)

So you'll take 25 μl from BSA stock and add it in your diluent (solvent). Since you're using Phosphate buffer as diluent, you'll add the BSA solution in it.

Your final volume should be 1 ml, and since volume taken from BSA stock solution is 25 μl, the rest should be the buffer.

So, 1000 - 25 = 975

You'll add 25 μl BSA stock in 975 μl of Phosphate buffer solution. This will make you final volume 1 ml.

Date PRE OD Stock solution - 2000 ug/me final volume of solution - 1 ml (Standard Formula - Mq Up į M₂ U2 lotto S OI 250 lua

LIIDUPIO Dato Pago dutor 1000 ug lime 1 M. U = M U le enote 2000 xu, - 1000x1000 ooo 2000 00 Lo q 2 500 ul from BSA stock @ 1

Date Page volume of BSA 50 ug/ml 4 25 ul t 250 ug/ml oool 125 ul t 500 ug/ml 25out 1000 ug/ml 500ul t 1250 ug/ml 625 ul t 150

I have added how much phosphate buffer solution you will need to make the final volume 1ml in every case  

Total volume of BSA which you'll be using in this is 2,275 μl.

Thank you